# Decoding Geometric Properties in Non-Random Data from First Information-Theoretic Principles

### v2

@@ 2.1 On the perturbation analysis of compressed data
While the above discussed cases study the bidimensional spaces, Fig. LABEL:bone below shows that our method can be scaled up to tridimensional spaces. As we will demonstrate in Section 2.2, we have found sufficient conditions for which the method remains sound for any multidimensional space. The challenges to tackle this generalization is presented in Section A.2.
2.2 Zero-knowledge one-way communication
The empirical results in this article suggest that AID is capable of harnessing (given the availability of adequate computational resources) some of the algorithmic-informational properties that we demonstrated in ideal and generalised scenarios. As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, most prominently for high algorithmically probable objects.
Our theoretical results show that the method employed to achieve the empirical findings in Section 2.1 is sound with regard to the infinite asymptotic limit when computational resources are unbounded; and it also shows that the method can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Our inference method is general enough to remain robust vis-a-vis any arbitrarily chosen alternative method to calculate the probabilities of the events, any arbitrarily chosen encoding-decoding scheme, programming language, or computation model, without the need of assuming some bias in the possible outcomes and prior beliefs.
Theorem 1 demonstrates that there are conditions for which decoding the message in the original multidimensional space $\mathcal{S}$ chosen from the one in which the message has the lowest algorithmic complexity (i.e., highest algorithmic probability, as assured by the universal distribution) becomes eventually more likely to be correct than decoding the message in a distinct multidimensional space $\mathcal{S}^{\prime}$, with or without noise. The underlying idea is that the larger the algorithmic complexity of the closer-to-randomness message candidate in the wrong multidimensional space with respect to the algorithmic complexity of the message in the correct multidimensional space, the larger the (algorithmic) likelihood [33, 34] of the correct multidimensional space with respect to the likelihood of the wrong closer-to-randomness multidimensional space. Thus, finding the multidimensional space that minimises the algorithmic complexity the most, eventually increases the probability of finding the correct one in comparison to the probability of finding the wrong closer-to-randomness one. Notice that this holds even if one assumes an arbitrarily different probability measure $\mathbf{P}$ as starting point.
We demonstrate that once sufficient conditions for the compressibility of objects, multidimensional spaces, and perturbations (e.g., noise) are met, one-way communication to the receiver becomes possible even if no prior knowledge about the encoding-decoding scheme chosen by the emitter is known. One can first simply look into the landscape of all possible partitions and low-complexity noise the receiver has access to while assuming no bias toward particular objects and partitions in this landscape. Then, despite having no knowledge about the encoding-decoding schemes and thus a lack of biases, one can infer with probability as high as one wishes that the message with the lowest algorithmic complexity is arbitrarily more likely to be correct than any other arbitrary computable method for assigning probabilities to message candidates.
One of the main conditions for this to hold is that the algorithmic complexity of the object $y$ (with or without noise) in the wrong multidimensional space should increase (i.e., move the message toward algorithmic randomness) sufficiently fast in comparison to the algorithmic complexity of the message in a correct multidimensional space. To this end, one can assume a message should contain a large amount of redundancies, and thus should be highly compressible with respect to its linear signal stream length $\left|y\right|$, such as the case in which $\mathbf{K}\left(y\middle|\mathcal{S}\right)=\mathbf{O}\left(\log\log\left(\left|y\right|\right)\right)$ holds, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity. In this manner, most low-complexity perturbations will transform the object and/or partition by sufficiently moving it toward algorithmic randomness so that the conditions are satisfied. This condition resembles the ones for the noisy-channel coding theorem in classical information theory [20], although our method can be understood as being more general than the classical version because: it holds for zero-knowledge communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26]. See also Section A.5 for further discussion on this and other mathematical phenomena. The proofs of all theorems can be found in Section 2.2.
3 Discussion
Our method is based on algorithmic information dynamics (AID) which combines tools from classical information theory and algorithmic information theory, in which algorithmic complexity values are approximated and combined with Shannon entropy by using the block decomposition method (BDM), in order to perform a perturbation analysis so that this algorithmic-information-based method can be used to avoid encoding biases in signal detection and processing. Our method is agnostic vis-a-vis endoding-decoding schemes and can be extended and used for bio- and technosignature detection and signal deconvolution. Such results relate information theory to fundamental areas of mathematics, such as geometry and topology, by means of compression and algorithmic probability.
As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM. In addition, although it may seem paradoxical at first glance—because one may initially think that the errors produced by the methods for approximating the algorithmic complexity would be inherited by the perturbation analysis phase itself—our theoretical and empirical results show that the perturbation analysis enables one to overcome the semi-computability limitation in algorithmic information theory, but only when, looking at the whole landscape of complexity values, one estimates with a sufficiently large amount of perturbations effected on the received signal. This is because the algorithmic complexities of the perturbed signals eventually become sufficiently larger than the complexity of the original message. Hence, a powerful enough method for approximating complexity, such as BDM, can eventually harness this divergence in order to enable the perturbation analysis to statistically distinguish the correct one (i.e., more compressible) from the wrong ones (i.e., more random). That is, an approximating method with enough computational resources can produce smaller errors than the difference of actual (although semi-computable) complexity values between the correct candidate and the perturbed ones, even though the exact complexity values remain ultimately unknown to the perturbation analyst (i.e., even though the analyst itself can never be sure it has enough computational resources in the first place). Further discussion of these and other topics can be found in Section A.5.
4 Methods
4.1 Formalism and basic concepts
@@
In this manner, one is always able to (uniquely) decode the received message into sufficient information to completely determine the multidimensional space configuration. One example of an encoding scheme for an arbitrary multidimensional space $\mathcal{S}$ consists in encoding it in the form of a companion tuple $\bm{\tau}$ as in [35, 36]. Finite and discrete multidimensional spaces with more intricate configurations other than whole number multiplication in the form of $m_{1}\times m_{2}\times\dots\times m_{k}$, where $m,k\in\mathbb{N}$, can, for example, be encoded by node-unaligned companion tuples $\bm{\tau}_{ua}$ as in [36, Definition 2]. Nevertheless, the theoretical framework of our method is agnostic vis-a-vis any arbitrarily chosen computational scheme for encoding multidimensional spaces.
Each partition of a $n$-bit-length linear signal stream corresponds to a distinct configuration of dimension lengths, and therefore to a distinct multidimensional space in which the total additive dimension lengths remains upper bounded by the constant $n$. For example, a picture of 4$\times$4 pixels can be partitioned into other dimension-length configurations leading to different non-squared shapes, like 2$\times$8, 8$\times$2, or 1$\times$16. Once a row (or, alternatively, a column) is fixed in a bidimensional space, the remaining rows and columns are dependent on the initial row fixation. In other words, each bidimensional partition is a single-variable dependency, which is the main rationale underlying the sweep in our algorithms employed to achieve the results in Section 2.1.
More formally, in the case of a (finite and discrete) bidimensional space $\mathcal{S}_{2}$, one may have different combinations of values $\left|{\mathcal{S}_{2}}_{x}\right|$ and $\left|{\mathcal{S}_{2}}_{y}\right|$ for each partition. However, for any partition one has it that both $1\leq\left|{\mathcal{S}_{2}}_{x}\right|\leq n$ and $\lfloor\frac{n}{\left|{\mathcal{S}_{2}}_{x}\right|}\rfloor=\left|{\mathcal{S}_{2}}_{y}\right|$ hold. Thus, because each partition is single-variable dependent on ${\mathcal{S}_{2}}_{x}$, for any $n$-bit-length linear signal stream one can encode each of its $2$D-partitions in $\mathbf{O}\left(\log\left(n\right)\right)$ bits. The same applies to $3$D-partitions (or to any partition from a finite number of dimensions), except for an extra number of bits upper bounded by a partition-independent constant.
@@
A (algorithmic) perturbation with low algorithmic complexity is one of these transformations where the algorithmic information content of the original object is mostly preserved under such perturbations. More formally, a low-complexity perturbation occurs when $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity, $x$ is the original encoded object, $y$ is the encoded object that results from the perturbation on $x$, and $\left|x\right|$ is the size of the object $x$.
Swapping all 0s for 1s, or swapping all black for white pixels in an image are examples of low-complexity perturbations (in these two cases in particular, one has it that $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(1\right)$). Randomly inserting or deleting a finite number of elements (such as creating or destroying edges in a graph or flipping a finite number of bits in a string) are also examples of low-complexity perturbations [26, 37, 24]. As we saw in the above paragraphs, changing the partition is also a low-complexity perturbation of the multidimensional space into which the object $x$ is embedded. This is because any new partition can always be encoded in $\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$ bits. Notice that a perturbation can change either the object (e.g., by flipping bits) or its multidimensional space (e.g., by reconfiguring the dimension lengths of a given partition to achieve another distinct partition) independently, or can change both at the same time. In any of these three options, except for the identity transformation, the encoded form of the message resulting from the perturbation will be distinct from the encoded form of the original message. Both changing the partition into which the object is embedded and introducing noise into the object itself (e.g., by flipping bits) are particular examples of (algorithmic) perturbations [26, 37, 24]. Notice that these properties of low-complexity perturbations underpin the methodology employed in Section 2.1.
@@
4.2 Toward a general theory for zero-knowledge one-way communication

### v3

@@ II-A On the perturbation analysis of compressed data
While the above discussed cases study the bidimensional spaces, Fig. LABEL:bone below shows that our method can be scaled up to tridimensional spaces. As we will present in Section III-B (and it is demonstrated in the Sup. Mat.), we have found sufficient conditions for which the method remains sound for any multidimensional space. Future research and challenges to tackle this generalization are fully explained in Appendices AE.
The next subsection presents novel empirical evidence of the suitability of not only structural perturbations, but also local bit-wise operations, as a robust strategy to find the best encode/decode scenario for a given message.
II-B Information content and bitwise perturbations
In introducing the bit-wise perturbation method put forward in this work and also illustrating the changes to compression, entropy and algorithmic complexity when considering different encoding methods, we analyse three types of strings:
  • A random binary string;
  • A string of random characters;
  • Real sentences from Darwin’s “Origin of Species”
The first step in the analysis is to bring all these strings into a common format. Thus, binary encodings were chosen such that both the strings of random characters and the sentences from Darwin’s book were binarised in a few different ways according to arbitrary UTF8-to-binary functions.
The proposed method is based on the premise that perturbations (with respect to both individual bits and spatial rearrangements of higher dimensional data) give us hints about the correct information content (including geometrical features) of a message. In order to assess the information content change, four metrics are measured for each perturbed string:
  • LZW: The length of the dictionary produced after applying a simple LZW lossless compression scheme to the input string;
  • Entropy: Shannon entropy value of the string (which could be Block Entropy for higher dimensional data);
  • BDM: The Block Decomposition Method (BDM) [31] is a strategy used to approximate the algorithmic or algorithmic complexity of objects like strings and graphs. Unlike conventional methods that focus on compressibility, such as lossless compression techniques, BDM divides a complex object into smaller, easier-to-handle components. These individual blocks are then evaluated based on their complexity using a reference database of small Turing machine outputs. This database provides a baseline for assessing the algorithmic content of each block. BDM builds upon the Coding Theorem Method (CTM), which relates algorithmic probability to algorithmic complexity by determining the chance that a random program could produce a given string on a universal Turing machine. By applying this logic to each block and summing their complexities, BDM estimates the complexity of the whole object. This approach is especially practical when calculating exact algorithmic complexity is too difficult due to the object’s size. By combining the complexity estimates of smaller blocks, BDM offers a broader, more refined view of complexity, surpassing the limitations of basic compressibility measures like Shannon entropy. This method reveals deeper insights into both structured and random data patterns.
  • z-lib: the length of the b64encode of the compressed zlib version of the string. This is an alternative to LZW lossless compression.
Thus, for our initial approach to the problem, the following sequence of steps will be applied:
  • Consider input string $str_{input}$;
  • Randomly shift “# of flipped bits” in the string, producing $str_{f,input}$;
  • Calculate the four metrics previously described for $str_{f,input}$;
  • Repeat the previous step 1024 times, keeping the same number of bits flipped but considering different bits each time.
In the following subsubsections the results for these perturbations are provided for each type of string. For the violin plots, the ticks inside each violin indicate the quartiles of the data, and the first and last results (no flipping or complete flipping) are represented as circles since the violin plot collapses to a single value distribution.
II-B1 Random binary strings
The first group of strings analysed is the one comprising random binary strings. So, let us consider the following random string of 402 binary numbers:
00011000111100011011010101010101000001001110101000010001101001010000001100011110001101101010101010100000100111010100001000110100101000 0001100011110001101101010101010100000100111010100001000110100101% 0000001100011110001101101010101010100000100111010100001000110100101000
01011000100101011010111001110101001000001100111011100000010101100100101100010010101101011100111010100100000110011101110000001010110010 0101100010010101101011100111010100100000110011101110000001010110% 0100101100010010101101011100111010100100000110011101110000001010110010
00010110100010000110010101000101010011010001101111100100100111110100001011010001000011001010100010101001101000110111110010010011111010 0001011010001000011001010100010101001101000110111110010010011111% 0100001011010001000011001010100010101001101000110111110010010011111010
11100010011111001011000111100000011010101110110111101111001010110111110001001111100101100011110000001101010111011011110111100101011011 1110001001111100101100011110000001101010111011011110111100101011% 0111110001001111100101100011110000001101010111011011110111100101011011
10011100000110011011000010111000000100001110011011111101101110101111001110000011001101100001011100000010000111001101111110110111010111 1001110000011001101100001011100000010000111001101111110110111010% 1111001110000011001101100001011100000010000111001101111110110111010111
01001010011111001100101010100010011001000101110111011111110001110110100101001111100110010101010001001100100010111011101111111000111011 0100101001111100110010101010001001100100010111011101111111000111% 0110100101001111100110010101010001001100100010111011101111111000111011
For such a string, there are exactly 201 0s and 201 1s. By running the simulations indicated, Fig. 2 was obtained. The analysis of Fig. 2 indicates that for binary random strings at maximal entropy, perturbations do not make the string “more random”, indicating that the information content does not change appreciably after bit flipping. This is a confirmation of the randomness of the input data.
Figure 2: Algorithmic perturbation analysis on strings of size 402. Random binary numbers: For a binary random string with balanced number of 0s and 1s, performing a bit flip for several bits (from 0 - original string, up to 402 - all the characters in the string) causes the resulting size of the LZW dictionary and the zlib encoded string to stay approximately the same (median value constant in the plots). The same is observed for Shannon entropy, which did not change appreciably and stayed close to its maximum value. This is to be expected since our original string is already in a maximal entropy state. On the other hand, BDM slightly decreases, which indicates that the bit flips brought the string to a slightly more informative state (less random too, because, in theory, perfect randomness requires the number of 1s and 0s to be the same, and this is not guaranteed for the flipped strings). Perturbing balanced binary random messages indicates that, overall, the only information gain possible in the process is obtained when the bit balance is broken, which is graphically represented as an up concavity curve for BDM values. This would be equivalent to tempering a coin and observing a series of its flips. Random characters sampled from a 95-character alphabet after vowel encoding: This new binary string is unbalanced, with 12.19 % of its bits being 1s. Bit flips were performed revealing that all the metrics tend to increase up to when half of the bits are flipped. This is to be expected since our original string, although randomly generated, considered an unbalanced encoding which increased information content. Besides, it is worth noting that the simple fact that the number of 0s and 1s is not the same provides information about the message. Therefore, the down concavity observed implies that the original message is more informative than any perturbation processed. This is because signals carrying meaning are far removed from randomness, and random perturbations make the text more random [33], even when the methods know nothing about words, grammar or anything linguistic.Random characters sampled from a 95-character alphabet after space encoding: This new binary string is highly unbalanced, with 1.49 % of its bits being 1s. Bit flips were performed revealing the same behaviour as observed for the vowel encoding. Compared to the vowel encoding, it is possible to see that all the metrics span for a larger interval when highly unbalanced encodings are considered.
II-B2 String of random characters
For this second class of strings, we will consider strings formed by randomly sampling a string pool. It is possible to consider the first 95 printable characters, for example, and by randomly sampling from this pool, obtain the following string with length 402:
\$eW*yOqM:Lqk#ifo1}o]‘404 bx+,QF3JGNh>~#E({z=5W!P3q%\$2rn |kZ4eLo33YCN1.O
[…]
jBt%m+Pq#tSfr_r\’3FFBy}z7dj)h?(Yi*/pX(i|)\’3@H)@vDvhu<2o&iD)y;F3\’Ivo5f
Zx+:fnYQw8>s) H‘bU<MoI/Uh^GGz&‘/jy(@Iz)mmH8CV%1U/9ZPOdc.s,">|V7[bd"%(A(
[…]
At the outset it is worth highlighting that even though the string was randomly built, it is not perfectly random since the frequency distribution of characters is not perfectly uniform. This is a direct consequence of the fact that we sampled our original 95-character alphabet 402 times (which means there are more samples than characters, implying oversampling of some of the alphabet elements - impossible to circumvent since 402/95 is not an integer). Therefore, this string does not have maximal entropy over the 95-character alphabet space. Besides, even if we could produce strings with maximal entropy over the 95-character space (for example, a string of size 380 with each character appearing exactly 4 times but in a random order), in general when converted to binary, the new converted string would not have maximal entropy anymore, since the encoding used will increase information content and diminish randomness. Fig. 3 presents a flowchart to illustrate this process.
Figure 3: Algorithmic perturbation flowchart for strings of characters randomly picked from a character pool. Even when randomly sampling from a character pool, it is possible that maximal entropy is not observed even in the original alphabet space. By using encoding to convert characters to binary, information content as well as order is introduced in the process, thereby making the converted string “less” random than the original one.
In general, when an encoding algorithm is considered, this implies conditioning the original random sampling, so that a maximal entropy string is almost always “organised” when converted to binary. It is interesting to understand how different encodings behave in that context.
  • Unbalanced encodings
By unbalanced we mean encodings which do not preserve the balance of 0s and 1s for each converted character. For example, let us consider the case where each vowel (upper and lower case - AaEeIiOoUu) is converted to 1 while other characters are converted to 0. This encoding is highly unbalanced since, on average, for long random strings taken from the pool, 10/95 of the converted characters will be 1s while 85/95 will be 0s. Fig. 2 presents the perturbation analysis for the binary string created after applying this encoding.
It is possible to repeat this same exercise with another, even more unbalanced encoding: each space is converted to 1 while all the other characters are converted to 0. For the string considered, this brings the percentage of 1s to a mere 1.49 %. The perturbation results are given in Fig. 2.
If we continue considering unbalanced encodings, it is natural to assess the performance of the traditional UTF-8 encoding. In that case, characters are represented as 8-bit strings, which are then concatenated to form the message to be analysed. Thus, for the initial 402 character message, the binary encoded string is of length 3216. The results for the perturbation analysis of the UTF-8 encoding of the random string of characters is presented in Fig. 4.
Figure 4: Algorithmic perturbation analysis on strings of size 3216. Random characters sampled from a 95-character alphabet after UTF-8 encoding: This binary string is almost balanced, with 47.42 % of its bits being 1s. Bit flips were performed for several bits (from 0 - original string, up to 3216 - all the characters in the string), revealing that with the exception of BDM, all the metrics only slightly change (less than 2%). BDM, on the other hand, points to a continuous increase up to when half of the bits are flipped. Once again, the down concavity observed implies that the original message is more informative than any perturbation processed. Compared to the other unbalanced encodings, as soon as a more balanced encoded string is seen, entropy and compression algorithms tend to lose their power to detect information content changes. Excerpt of Darwin’s “Origin of Species” after UTF-8 (unbalanced) encoding: This binary string is almost balanced, with 45.55 % of its bits being 1s. Bit flips were performed revealing that BDM and z-lib were the most sensitive metrics to information content changes. The down concavity observed implies that the original message is more informative than any perturbation processed.
Based on all the examples presented for unbalanced encodings, it is evident that even randomly generated strings may not present maximal entropy when converted to binary. As a matter of fact, as soon as the encoded binary strings start to approach a balance of 0s and 1s, entropy and compression algorithms lose their power to detect information content changes. BDM, on the other hand, presented good results even in more challenging scenarios. Also, if we apply the method to any string and see that the original string (without flipping) is the one with minimum metrics, then the method could not improve on our initial guess about the information content of the message (the message is what it is). This may be true for some 1D strings. On the other hand, for multidimensional cases, there will be downward spikes which will indicate a “more suitable” configuration, especially because we may not start the perturbation from the “correct” string as we did in the examples above.
  • Balanced encodings
In contrast, it is possible to state that a given encoding is balanced if the number of 0s and 1s is balanced (equal) for every character converted. This can be achieved by simply concatenating the UTF-8 encoding of a given character (8-bits) with its bitwise inverted encoding (additional 8-bits), producing a 16bit encoding per character. Such an encoding will always have the same number of 1s and 0s. In this case, the original 402 characters become 6432 bits and the results are presented in Fig. 5.
Figure 5: Algorithmic perturbation analysis on strings of size 6432. Random characters sampled from a 95-character alphabet after balanced encoding: This binary string is perfectly balanced, with 50 % of its bits being 1s. Bit flips were performed for several bits (from 0 - original string, up to 6432 - all the characters in the string), revealing that except from BDM and z-lib, all the metrics present small changes (less than 2%). BDM and z-lib, on the other hand, present a continuous increase up to when half of the bits are flipped. BDM is more sensitive than z-lib, as the size of the plateau is smaller with longer lateral climbs. Once more, the down concavity observed implies that the original message is more informative than any perturbation processed. Overall, entropy and LZW lose their power as proxies of information content whenever balance is observed for the binary string. Excerpt of Darwin’s “Origin of Species” after balanced encoding: This binary string is also perfectly balanced, with 50 % of its bits being 1s. Bit flips were performed, once again revealing the capabilities of BDM and z-lib to assess the information content of the message, and general analysis implies that the original message is more informative than any perturbation processed.
II-C Sentences from Darwin’s “Origin of Species”
Following the same rationale, we shall apply the perturbation techniques to encoded binary strings obtained after applying both unbalanced (strict UTF-8) and balanced (UTF-8 concatenated with its bitwise inverted encoding) encodings to the following sentences from Darwin’s “Origin of Species” (total length of 402 characters):
we have many slight differences which may be called individual differences, such as are known frequently to appear in the offspring from the same parents, or which may be presumed to have thus arisen, from being frequently observed in the individuals of the same species inhabiting the same confined locality. No one supposes that all the individuals of the same species are cast in the very same mould
The perturbation analyses for unbalanced and balanced encodings of the excerpt from Darwin’s “Origin of Species” are found in Figs. 4 and 5, respectively.
Figs. 4 and 5 reveal that the results observed for random strings are also true for “real” sentences: BDM and z-lib are sensitive to information content change and the simpler LZW algorithm together with entropy rapidly lose their abilities to detect information content change whenever the balance of 1s and 0s in the encoded binary string is approached.
The results in this section demonstrate that most single-bit perturbations in a sequence that encodes a message result in a more random-looking sequence. These results vary depending on what mapping we choose to binarise the sequences, but in general, the more compressible the object is, the more a perturbation increases its randomness across various mappings (we also explored other arbitrary binarisation functions, with similar results). (See also Sup. Mat. and [33] for mathematical proofs of such a phenomenon). In other words, these examples show that sequences that carry meaningful content are quantitatively different according to various measures. Any deviation (i.e., perturbation) from the encoded sequences causes randomness to increase. Confirming our theoretical results in this paper (see Section III-B and Sup. Mat.), our results suggest that an incoming signal with low randomness is more likely to encode a message than one that has more randomness.
Fig. 6 demonstrates that the same phenomenon is observed in audio signals, where the perturbation approach has been applied to an audio recording of the words spoken from the Apollo 13 mission: “Houston, we’ve had a problem”. This message was transmitted on April 13, 1970 at a sample rate frequency of 11025 Hz on a single channel. When compressed, the message is 106 250 bytes and 106 320 characters long. Here, the message was scrambled several times. The histograms on the bottom left show the difference between compression lengths in bytes and string lengths (length of zlib after b64encoding) of the scrambled messages with respect to the compressed results of the undisturbed message. Each of these scramblings results in an increase in randomness. Additionally, if only the beginning and end of each word is perturbed, the resulting randomness increases smoothly with each perturbation, as shown in the bottom right plot. These results indicate that perturbations of the original message, including scrambling elements of the original message in a different order, will result in a message of higher complexity and thus greater randomness. As a result, the original encoding of the message (the one with an interpretable meaning) is the one with the lowest complexity.
Figure 6: Top: Sound waves of the words spoken from the Apollo 13 mission: “Houston, we’ve had a problem.”. The message was transmitted on April 13, 1970 at a sample rate frequency of 11025 Hz on a single channel. Bottom left: Histograms of the change in compression lengths of scrambled versions of the same message from the length of the original compressed message. By change we mean byte counts minus 106 250 bytes and string lengths minus 106 320. Bottom right: Small perturbations of the original message by scrambling only the beginning and end of each word shows a smooth transition from low to high randomness. The lowest complexity signal indicates the correct (original) signal. The file was processed in FLAC format (Free Lossless Audio Codec) from a lossless file, with no audio data discarded.
Having investigated how bitwise perturbations could point towards better candidates for intelligible messages, it is also interesting to see how rearranging the message could potentially indicate suitable candidates for their respective multidimensional space, such as their natural geometrical features. This was preliminarily explored in [30] and shall be further investigated in the next subsection.
II-D Information content and structural perturbations
All the examples presented so far were created based on some sort of encoding from a high information content message to binary. Thus, in general, it would be expected that groups of binary elements in the string, not single bits, would carry the actual information content. In this regard, it is interesting to reshape the original binary string into several different configurations, then measure some of the metrics considered during the bitwise perturbations. Considering that z-lib and LZW algorithms are both compression algorithms, we chose to evaluate only z-lib, BDM and entropy.
Since the size of the original binary message of size $s$ may not necessarily be reshaped as $m\times n$ (i.e., the original size is not the product of $m$ and $n$), it is important to set the maximum information loss allowed in this reshaping process. This can be set, for example, as a 1% loss for the purposes of our investigation, which indicates that we are looking for $m$ and $n$ such that $mn\geq 0.99s$, where $s$ is the length of the original encoded binary message. Also, there may be more than one pair of $m$ and $n$ values which satisfy such an inequality. Thus, we set $n$ as the maximum second dimension for a given $m$ value. So, the dimensions actually searched are $(m,n)$, both natural numbers, such that $n=max(n_{1},n_{2},...n_{j})$, where $mn_{i}\geq 0.99s$, for $i=1,...,j$.
Besides, we will compare values for which small information loss is possible (since $|s-mn|\leq 0.01$), making it important to take this into account while calculating the metrics. The following changes are then considered:
  • the BDM value is divided by the number of 4x4 grids used to make the partitioning of the original 2D array (i.e. $(m-3)(n-3)$);
  • Entropy value is changed to Block Entropy, where the block size is equal to $n$;
  • zlib b64encoding length is divided by the ratio of bits kept (i.e. $mn/s$).
To make the comparisons more clear, for every metric, a MinMaxScaler was applied, such that the results range from 0 to 1.
For the excerpts from Darwin’s book, the results of this reshaping process are presented in Fig. 7. This figure reveals that structural perturbations are able to suggest possible structural arrangements for the message in the space into which it is embedded, allowing one to shortlist a few candidate dimensions based on downward peaks of BDM in the low-BDM region. This is done in a totally agnostic fashion. Since this method showed promising results for text and audio data, it would now be of interest to explore images.
Figure 7: Left: Structural perturbations on the UTF-8 binary encoding of Darwin’s book excerpt. Right: Structural perturbations on the balanced encoding of Darwin’s book excerpt. For both cases, the ticks on the radar plot are the values of $m$ obtained after reshaping the initial string as a $m\times n$ array. Also, the “correct” shape of the message is represented in red and is observed for downward BDM peaks in the low-BDM region. It can be seen that the balanced encoding boosted the detection of the original UTF-8 encoding (which would be equivalent to a $804\times 8$ array), since the algorithmic complexity of the bit shifting operation is small.
So far, robust empirical evidence has been presented to support the new method hereby proposed. In the next section, theoretical foundations of the method shall be presented.
III Methods
III-A Formalism and basic concepts
@@
In this manner, one is always able to (uniquely) decode the received message into sufficient information to completely determine the multidimensional space configuration. One example of an encoding scheme for an arbitrary multidimensional space $\mathcal{S}$ consists in encoding it in the form of a companion tuple $\bm{\tau}$ as in [34, 35]. Finite and discrete multidimensional spaces with more intricate configurations other than whole number multiplication in the form of $m_{1}\times m_{2}\times\dots\times m_{k}$, where $m,k\in\mathbb{N}$, can, for example, be encoded by node-unaligned companion tuples $\bm{\tau}_{ua}$ as in [35, Definition 2]. Nevertheless, the theoretical framework of our method is agnostic vis-a-vis any arbitrarily chosen computational scheme for encoding multidimensional spaces.
Each partition of a $n$-bit-length linear signal stream corresponds to a distinct configuration of dimension lengths, and therefore to a distinct multidimensional space in which the total additive dimension lengths remains upper bounded by the constant $n$. For example, a picture of 4$\times$4 pixels can be partitioned into other dimension-length configurations leading to different non-squared shapes, like 2$\times$8, 8$\times$2, or 1$\times$16. Once a row (or, alternatively, a column) is fixed in a bidimensional space, the remaining rows and columns are dependent on the initial row fixation. In other words, each bidimensional partition is a single-variable dependency, which is the main rationale underlying the sweep in our algorithms employed to achieve the results discussed in Section II.
More formally, in the case of a (finite and discrete) bidimensional space $\mathcal{S}_{2}$, one may have different combinations of values $\left|{\mathcal{S}_{2}}_{x}\right|$ and $\left|{\mathcal{S}_{2}}_{y}\right|$ for each partition. However, for any partition one has it that both $1\leq\left|{\mathcal{S}_{2}}_{x}\right|\leq n$ and $\lfloor\frac{n}{\left|{\mathcal{S}_{2}}_{x}\right|}\rfloor=\left|{\mathcal{S}_{2}}_{y}\right|$ hold. Thus, because each partition is single-variable dependent on ${\mathcal{S}_{2}}_{x}$, for any $n$-bit-length linear signal stream one can encode each of its $2$D-partitions in $\mathbf{O}\left(\log\left(n\right)\right)$ bits. The same applies to $3$D-partitions (or to any partition from a finite number of dimensions), except for an extra number of bits upper bounded by a partition-independent constant. Notice that one can restrict the search for partitions for which the resulting product of $\left|{\mathcal{S}_{2}}_{x}\right|$ and $\left|{\mathcal{S}_{2}}_{y}\right|$ are as close as possible to the original value of $n$. For example, this is employed in Section II-D with a constant loss rate, thereby not changing the resulting $\mathbf{O}\left(\log\left(n\right)\right)$ bits necessary to encode a particular partition in such a restricted search.
@@
A (-n algorithmic) perturbation with low algorithmic complexity is one of these transformations where the algorithmic information content of the original object is mostly preserved under such perturbations. More formally, a low-complexity perturbation occurs when $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity, $x$ is the original encoded object, $y$ is the encoded object that results from the perturbation on $x$, and $\left|x\right|$ is the size of the object $x$.
Swapping all 0s for 1s, or swapping all black for white pixels in an image are examples of low-complexity perturbations (in these two cases in particular, one has it that $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(1\right)$). Randomly inserting or deleting a finite number of elements (such as creating or destroying edges in a graph or flipping a finite number of bits in a string) are also examples of low-complexity perturbations [26, 33, 24]. As we saw in the above paragraphs, changing the partition is also a low-complexity perturbation of the multidimensional space into which the object $x$ is embedded. This is because any new partition can always be encoded in $\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$ bits. Notice that a perturbation can change:
  • either the object (e.g., by flipping bits), thus called an object perturbation or a content perturbation;
  • or can change its multidimensional space (e.g., by reconfiguring the dimension lengths of a given partition to achieve another distinct partition) independently, thus called a structural perturbation or context perturbation;
  • or can change both at the same time.
In any of these three options, except for the identity transformation, the encoded form of the message resulting from the perturbation will be distinct from the encoded form of the original message. Both changing the partition into which the object is embedded and introducing noise into the object itself (e.g., by flipping bits) are particular examples of (algorithmic) perturbations [26, 33, 24]. Notice that these properties of low-complexity perturbations underpin the methodology employed in Section II-A and Sup. Mat..
@@
III-B Theoretical results